OpenStudy (anonymous):

Evaluate the expression below when x = -4. 4x2 - 2x × 15

OpenStudy (lalaly):

\[4(-4)^2-2(-4) \times 15\]=\[64+8 \times 15\]=

OpenStudy (anonymous):

Plug in x = -4 4x^2 -2x X 15 4(-4)^2 -2(4) X 15 4*16 - 8X15 64 - 120 -56

myininaya (myininaya):

4(-4)^2-2(-4)*15 4(16)-2(-4)(15) 4(16)+2(4)(15) 4(16)+8(15) 64+120 184

OpenStudy (lalaly):

lols i forgot the 184